If one elevator is on the 139th floor and travels two seconds per floor, and the other elevator is on the 14th floor and travels three seconds per floor, on what floor will they meet. What formula would you use to solve this problem. thank you.
Answer by richwmiller(17219)

139-14=125 floors difference
u+d=125,
3u=2d
d=75
u=50
50+14=64 50 floors going up
139-75=64 75 floors going down
3*50=150 seconds going up
2*75=150 seconds going down